# MMSci: A Dataset for Graduate-Level Multi-Discipline Multimodal Scientific Understanding

### v2

@@ 5 Training Resources
Task-specific Multimodal Training Data
We created a task-specific multimodal training dataset for visual supervised fine-tuning, formatted as single-turn or multi-turn conversations focused on figure captions. The two tasks used for benchmark evaluation, multiple-choice questions and figure captioning, are structured as single-turn interactions. For figure captioning, only abstract-grounded data is included. Additionally, we transformed figure captions into multi-turn conversations, where a human asks about content in different sub-figures, and the assistant responds with the corresponding sub-caption in each turn. To introduce diversity, we generated various conversation templates using GPT-4 to vary human instructions. All model responses are sourced from the original articles to ensure data quality. This process produced 108,843 multi-turn conversations, resulting in a training dataset with over 1 million instances.
Table 3: Performance on scientific figure captioning. Abs. denotes whether grounded on abstract for captioning. The LLM-based evaluation results, using GPT-4o, are reported on a randomly selected subset of 100 samples. The best results are highlighted in bold, with the second-best underlined.
Model Abs. Overlap-based Similarity-based LLM-based
Rouge1 Rouge2 RougeL Meteor BertScore FActScore* G-Eval
@@
Kosmos2 20.69 2.07 11.69 14.53 77.51 0.97 1.03
LLaVA1.5-7B 19.25 1.96 12.56 11.80 79.93 4.39 1.12
LLaVA1.6-Mistral-7B 18.48 2.83 10.97 20.45 79.53 5.57 1.18
Qwen-VL-7B-Chat 23.44 4.02 14.78 15.34 81.95 3.25 1.27
InternVL2-2B 15.12 2.69 9.60 17.74 78.89 6.74 1.78
InternVL2-8B 18.89 3.83 11.39 21.07 79.41 8.54 2.60
IDEFICS2-8B 12.83 1.93 9.40 6.51 80.30 2.78 1.36
IDEFICS3-8B-Llama3 15.51 2.89 10.11 19.09 78.65 7.52 1.68
MiniCPM-V-2.6 28.38 5.15 14.57 24.84 81.19 11.70 2.94
Llama3.2-11B-Vision 19.93 4.32 12.98 21.21 78.89 8.49 2.47
Qwen2-VL-2B 19.80 4.31 12.74 21.39 80.03 10.36 2.35
Qwen2-VL-7B 21.42 4.92 12.96 23.88 80.06 10.28 3.46
Qwen2-VL-2B-MMSci 29.28 8.42 19.77 19.75 83.56 9.19 2.73
@@
Gemini-1.5-Flash 25.39 6.77 15.49 26.82 81.10 8.31 3.71
Gemini-1.5-Pro 28.59 7.62 16.38 27.06 81.13 14.70 3.77
Claude-3.5-Sonnet 27.68 5.57 15.54 26.32 81.76 9.55 3.57
GPT-4V 27.37 6.02 14.86 26.62 81.75 14.35 3.71
GPT-4o 27.53 6.82 15.59 27.02 81.11 13.40 4.03
@@
Kosmos2 23.68 3.59 11.81 19.54 79.09 4.18 1.40
LLaVA1.5-7B 23.16 3.53 13.97 14.54 81.20 9.29 2.01
LLaVA1.6-Mistral-7B 21.52 4.13 12.70 21.49 80.84 7.85 1.45
Qwen-VL-7B-Chat 25.49 4.47 15.55 16.02 81.87 9.34 1.62
InternVL2-2B 19.19 3.89 11.74 18.45 80.88 10.52 2.13
InternVL2-8B 20.51 4.85 12.30 22.66 80.57 10.15 2.99
IDEFICS2-8B 15.27 2.62 10.81 8.06 80.30 5.31 1.94
IDEFICS3-8B-Llama3 17.99 3.66 11.28 20.62 79.42 8.05 1.97
MiniCPM-V-2.6 30.41 5.83 15.36 25.09 82.68 13.65 2.95
Llama3.2-11B-Vision 18.16 4.47 11.24 22.63 79.63 9.97 2.16
Qwen2-VL-2B 23.08 5.18 14.47 21.77 81.23 12.02 2.62
Qwen2-VL-7B 24.19 6.16 14.45 26.00 81.21 10.57 3.43
Qwen2-VL-2B-MMSci 30.80 9.29 20.70 21.44 83.78 14.18 3.18
@@
Gemini-1.5-Flash 26.74 7.47 16.03 28.71 81.80 10.27 4.08
Gemini-1.5-Pro 28.71 7.73 16.89 28.91 81.93 13.98 4.11
Claude-3.5-Sonnet 29.60 6.71 16.65 27.52 81.76 12.30 4.03
GPT-4V 28.45 7.01 15.65 27.62 82.37 19.64 4.15
GPT-4o 28.85 7.79 16.36 28.37 81.84 19.11 4.25
Interleaved Text and Image Data for Pre-training
MMSci includes full article content and figures, naturally forming interleaved text and image data suitable for pre-training MLLMs (Lin et al., 2023). We discuss the utilization of this interleaved data in Section 7.
6 Benchmark Evaluation Results
Evaluated Models
We evaluated a range of open-source and proprietary MLLMs, including Kosmos-2 (Peng et al., 2023), Qwen-VL-7B-Chat (Bai et al., 2023a), Qwen2-VL-2B, and Qwen2-VL-7B (Wang et al., 2024a), the LLaVA1.5 and LLaVA-NeXT(1.6) models (Liu et al., 2024; 2023a), IDEFICS2 (Laurençon et al., 2024b) and IDEFICS3 (Laurençon et al., 2024a), the InternVL2 series Chen et al. (2024a), and Llama3.2-11B-Vision (Team, 2024). For proprietary models, we evaluated Gemini-1.5-Flash and Gemini-1.5-Pro (Reid et al., 2024), Claude-3-Opus (Anthropic, 2024a), Claude-3.5-Sonnet (Anthropic, 2024b), GPT-4V, and GPT-4o (Achiam et al., 2023). The exact model versions used are detailed in Appendix A.2.1. Moreover, we fine-tuned a Qwen2-VL-2B model using our task-specific training data for one epoch, resulting in Qwen2-VL-2B-MMSci.
Table 4: Accuracies (%) and ranks of different models on multiple-choice questions.
Model Fig2Cap SubFig2Cap SubCap2Fig Overall
Accuracy Rank Accuracy Rank Accuracy Rank Accuracy Rank
Open-source Models
Kosmos2 23.99 19 23.95 17 24.33 16 24.09 19
LLaVA1.5-7B 32.74 18 24.31 16 22.80 18 26.75 17
LLaVA1.6-Mistral-7B 34.76 17 20.38 18 24.15 17 26.60 18
Qwen-VL-7B-Chat 39.56 16 19.93 19 27.83 15 29.23 16
InternVL2-2B 42.76 15 33.07 13 38.42 13 38.18 13
InternVL2-8B 52.78 11 49.60 10 40.13 12 47.62 12
IDEFICES2-8B 48.65 14 25.83 15 21.10 19 32.21 15
IDEFICES3-8B-Llama3 50.42 13 28.43 14 29.98 14 36.57 14
MiniCPM-V-2.6 53.20 10 58.27 8 61.67 9 57.61 8
Llama3.2-11B-Vision 54.97 9 45.04 11 71.18 7 57.00 9
Qwen2-VL-2B 60.61 7 37.62 12 55.12 11 51.30 11
Qwen2-VL-7B 66.16 4 73.10 7 79.80 4 72.87 5
Qwen2-VL-2B-MMSci 78.62 1 83.02 3 83.57 3 81.67 1
Proprietary Models
Gemini-1.5-Flash 54.77 6 77.84 5 64.41 8 65.24 7
Gemini-1.5-Pro 62.79 5 81.41 4 77.16 5 73.52 4
Claude-3-Opus 52.19 12 53.17 9 60.23 10 55.13 10
Claude-3.5-Sonnet 68.77 2 85.34 2 87.16 1 80.18 2
GPT-4V 60.43 8 75.07 6 76.12 6 70.45 6
GPT-4o 67.42 3 87.40 1 84.65 2 79.57 3
Human (CS Graduates) 49.67 (±5.56$\pm$ 5.56)± 5.56 ) 69.00 (±$\pm 4.97$) 59.33 (±$\pm 5.44$) 58.45 (±$\pm 2.73$)
Random Guess 25.86 (±$\pm 0.94$) 24.63 (±$\pm 0.70$) 20.62 (±$\pm 0.73$) 23.24 (±$\pm 0.49$)
Scientific Figure Captioning Results
@@ 6 Benchmark Evaluation Results
The captioning results are shown in Table 3. We observe that grounding captions in the article abstract consistently improves generation quality by providing essential contextual background. This suggests that models struggle to generate captions that fully match the ground truth’s nuanced semantics, format, and style. Only our fine-tuned model seems to have learned these subtle semantic and stylistic details from the training data.
In terms of LLM-based metrics, which evaluate quality beyond semantic nuance, open-source models still significantly underperform compared to proprietary models, revealing even greater deficiencies. While MiniCPM-V-2.6 and Qwen2-VL-7B come close to proprietary models in FActScore, which measures precision, their G-Eval scores are notably lower. This suggests that these models fail to capture as much detail across entire figures as proprietary models do. Even our fine-tuned model does not achieve satisfactory G-Eval scores, despite comparable precision in FActScore, underscoring the high demands of this task on model capability. Although proprietary models, particularly GPT-4o and GPT-4V, achieve strong G-Eval scores, this may be because GPT-4o was used as the evaluator. When considering FActScore, all models still fall short of reasonable performance, highlighting their deficiency in precisely describing scientific figures and leaving room for improvement.
Multi-choice Question Results
The results of the multiple-choice questions are shown in Table 4. We observe a significant performance gap across different models and settings. The figure-to-caption (Fig2Cap) task is the most challenging, requiring models to summarize entire multi-panel figures and distinguish the correct summary from several similar ones within the same paper (examples can be found in Figure 8 in the Appendix). Our fine-tuned model achieved the best performance, outperforming the strongest proprietary model by around 10%. Notably, the gap between open-source and proprietary models was smallest in this setting. Open-source models, such as Qwen2-VL-2B and Qwen2-VL-7B, performed comparably to proprietary models. However, in the SubFig2Cap and SubCap2Fig tasks, the performance of open-source models lagged significantly behind proprietary models, suggesting a limitation in their ability to identify nuanced content within figures.
Overall, some open-source models, Kosmos2, LLaVA1.5, LLaVA1.6, Qwen-VL-7B-Chat, InternVL2-1B, and IDEFICS2-8B, performed no better than random guessing. In contrast, others, including MiniCPM-V-2.6, Llama3.2-11B-Vision, and Qwen2-VL-7B, demonstrated strong competitiveness with proprietary models. Interestingly, the proprietary model Claude-3-Opus performed significantly worse. The top-performing proprietary models were Claude-3.5-Sonnet and GPT-4o, but our fine-tuned Qwen2-VL-2B-MMSci achieved the highest overall performance. This suggests that the gap between proprietary and open-source MLLMs is narrowing, and that a smaller 2B model fine-tuned on our dataset can effectively bridge this gap, highlighting potential for further improvements in these models. When considering the results of scientific figure captioning, Claude-3.5-Sonnet and GPT-4o emerged as the top performers.
We asked three CS master students to evaluate our benchmark. Specifically, we sampled 1-2 questions from each subject, resulting in a subset of 100 questions per setting. The results are shown in Table 4. The CS graduates achieved an overall accuracy of 58.45, lower than that of some MLLMs. This outcome is understandable, given their limited knowledge in many of the scientific fields covered. It highlights the challenge of our task, which demands graduate-level expertise in specific disciplines.
Figure 4: Accuracies of the top 10 performing models across subjects in three major categories under the Fig2Cap setting.
Performance Breakdown Across Subjects
Figure 4 shows the performance of 10 top-performing models across three of the five major categories. The physical and biological sciences are generally more challenging, while performance in health science subjects varies significantly. Some subjects are difficult for all models, with top performance falling below 50%, like Rheumatology. Our fine-tuned model leads in the most subjects overall. Additionally, Claude-3.5-Sonnet excels in ecology, microbiology, psychology, and molecular medicine; Gemini-1.5-Pro leads in engineering, genetics, and biochemistry; GPT-4o performs best in evolution and pathogenesis; and Qwen2-VL-7B leads in mathematics and computing.
@@ 7 A Case Study in Material Sciences
Material science is the subject with the most articles and figures in our dataset. It is an important and highly interdisciplinary field, requiring knowledge from various subjects. Therefore, we conducted a case study to enhance material science knowledge using our dataset. There has been research on using language models for material science tasks (Walker et al., 2021; Rubungo et al., 2023; Miret & Krishnan, 2024). A recent study (Gruver et al., 2024) achieved promising results by utilizing LLaMA2 (Touvron et al., 2023) for material generation. In this study, material crystal structures were represented as text strings, and the LLaMA2 model was trained to generate these structure strings. However, LLaMA2 may lack sufficient scientific knowledge to fully comprehend the principles of material generation. Therefore, we explored the continuous pre-training of LLaMA2 using our interleaved scientific article and figure data, aiming to enhance the model’s performance on the stable material generation task.
Visual Pre-Training on MMSci
We continuously pre-trained the LLaMA2-7B model on our collected interleaved article text and figure images, using data within materials science as well as other eight related subjects in the same Physical Science category. To achieve that, we leverage LLaVA’s architecture (Liu et al., 2024), equipping LLaMA2 with a pre-trained CLIP ViT-L/14-336 (Radford et al., 2021) as the visual encoder and a 2-layer MLP as the projector. During training, we initially kept the LLM frozen and used data from general domains provided by (Liu et al., 2024) to initialize the projector. We then trained the model on the interleaved text and image data from general domains in MMC4 (Zhu et al., 2024) to further develop its image perception abilities, followed by our collected interleaved articles and figures in MMSci to infuse scientific knowledge. In this stage, we tuned both the LLM and the projector, for one epoch. For the resulting multimodal model, we use its LLM part, named LLaMA2-7B-MMSci, for the subsequent material generation.
@@
Fine-tuning for Materials Generation
Given the LLM, we further fine-tune it for the material generation task as in (Gruver et al., 2024). Specifically, periodic materials are characterized by a unit cell that repeats infinitely in all three dimensions. Each unit cell is specified by its side lengths ($l_{1}$, $l_{2}$, $l_{3}$) and angles ($\theta_{1}$, $\theta_{2}$, $\theta_{3}$). Within this lattice structure, there are $N$ atoms, each identified by an element symbol, $e_{i}$, and a set of 3D coordinates ($x_{i}$, $y_{i}$, $z_{i}$). Tzhe structure of a bulk material $C$ can be represented by:
@@
Consistent with prior work (Xie et al., 2021; Gruver et al., 2024), we experiment on the MP-20 dataset (Jain et al., 2013), which contains 45,231 stable materials. Therefore, an effective generative model trained on MP-20 is expected to generate new crystals that are at lease metastable. We construct the training data from these materials with two types of prompts: conditional generation (with one or multiple conditions) and infilling prompts, where partial crystal structure strings are masked and the model generates the masked parts. We train the model for one epoch, as training for more epochs reduces the diversity and coverage of generated materials.
@@
Gruver et al. (2024)
@@
Ours
@@
Results
We evaluate unconditional generation where no conditions are provided, allowing the model to generate potential stable materials for discovery. Consistent with (Xie et al., 2021; Gruver et al., 2024), we sample 10,000 generations with a temperature of 0.7. The evaluation focuses on four key aspects: validity, which ensures adherence to physical constraints; coverage and property metrics, which measure the alignment between the ground truth and the sampling distribution; and stability checks, which determine the percentage of samples deemed metastable by M3GNet (Chen & Ong, 2022) and stable by DFT (Hafner, 2008). As observed in Table 5, the LLaMA2-7B model, after being continuously pre-trained on our interleaved articles and figures and multi-task fine-tuning, consistently yields good results and achieves the best compositional validity, coverage precision, metastability, and stability. This underscores the benefit of our data in enhancing the generative model’s acquisition of scientific knowledge.
@@ Results
Ablation Studies
To understand the factors contributing to LLaMA2-7B-MMSci’s performance, we explored different pre-training data configurations: using only interleaved data from either MMC4 (general interleaved data) or MMSci, using interleaved data from MMC4 combined with text-only data from MMSci, and using no additional pre-training data, followed by the same fine-tuning setup. As shown in Figure 6, the text-only and interleaved data from MMSci achieved the top-2 overall performance when combined with MMC4 which equips the model to effectively read text and interpret images within scientific articles. Using both articles and figures led to better performance than using text-only data from MMSci, highlighting the importance of understanding both figures and content in scientific literature. In contrast, using only general domain data from MMC4 did not result in improvements, and directly training on MMSci even slightly decreased performance in structure validity. This is likely because incorporating visual information can confuse the model if it has not been sufficiently pre-trained with general interleaved data. Overall, continuous pre-training on our data shows the potential to infuse scientific knowledge that enhances downstream tasks.

### v3

@@ 5 Training Resources
Task-specific Multimodal Training Data.  We created a multimodal training dataset for visual fine-tuning, comprising both single-turn interactions (multiple-choice questions and abstract-grounded figure captioning) and multi-turn conversations about figure interpretation. The multi-turn conversations were generated by transforming figure captions into question-answer pairs about specific sub-figures, with diverse conversation templates generated by GPT-4. All responses were derived from original article content to ensure data quality. This approach yielded over 1 million training instances across 108,843 conversations. Fine-tuning Qwen2-VL-7B (Wang et al., 2024a) on this dataset for one epoch resulted in Qwen2-VL-7B-MMSci.
Table 3: Performance on scientific figure captioning. B2, RL, M, BS, CD, FS, and GE denote BLEU-2, ROUGE-L, METEOR, BERTScore, CIDEr, FActScore, and G-Eval, respectively. *The LLM-based evaluation results, using GPT-4o, are reported on a randomly selected subset of 200 samples. The best results are bolded, with the second-best underlined.
Model Image-only Captioning Abstract-grounded Captioning
B2 RL M BS CD FS* GE* B2 RL M BS CD FS* GE*
@@
Kosmos2 4.94 11.69 14.53 77.51 0.97 0.87 1.12 2.90 11.81 19.54 79.09 1.62 3.99 1.39
LLaVA1.5-7B 3.15 12.56 11.80 79.93 0.17 3.89 1.08 3.70 13.97 14.54 81.20 0.76 9.07 2.02
LLaVA1.6-Mistral-7B 2.8 10.97 20.45 79.53 0.08 5.17 1.23 3.90 12.70 21.49 80.84 0.48 7.67 1.47
Qwen-VL-7B-Chat 10.02 14.78 15.34 81.95 1.43 3.06 1.28 8.80 15.55 16.02 81.87 2.78 9.14 1.64
InternVL2-2B 1.69 9.60 17.74 78.89 0.03 5.99 1.76 2.27 11.74 18.45 80.88 0.96 10.38 2.17
InternVL2-8B 2.50 11.39 21.07 79.41 0.00 8.01 2.63 3.74 12.30 22.66 80.57 0.02 9.98 3.00
InternVL2-26B 4.18 13.26 24.21 81.02 0.19 12.43 3.01 5.21 14.92 23.19 80.27 2.30 12.31 3.20
IDEFICS2-8B 6.18 9.40 6.51 80.30 0.21 2.56 1.40 6.96 10.81 8.06 80.30 0.65 5.17 1.96
IDEFICS3-8B-Llama3 1.85 10.11 19.09 78.65 0.00 7.26 1.71 2.33 11.28 20.61 79.42 0.15 7.71 1.98
MiniCPM-V-2.6 4.75 14.57 24.84 81.19 1.42 11.15 2.96 6.11 15.36 25.09 82.68 3.27 12.93 2.95
Llama3.2-11B-Vision 2.68 12.98 21.21 78.89 0.08 8.27 2.46 2.60 11.24 22.63 79.63 0.00 9.55 2.18
Qwen2-VL-2B 3.45 12.74 21.39 80.03 0.38 9.94 2.31 5.68 14.47 21.77 81.23 1.43 11.88 2.64
Qwen2-VL-7B 3.60 12.96 23.88 80.06 0.00 10.03 3.39 4.73 14.45 26.00 81.21 0.19 10.36 3.45
Qwen2-VL-7B-MMSci 19.13 21.78 20.89 84.33 3.73 18.17 3.21 21.77 23.46 23.23 84.92 5.94 20.59 3.54
@@
Gemini-1.5-Flash 4.84 15.49 26.82 81.10 0.08 8.18 3.70 5.24 16.03 28.71 81.80 0.00 10.14 4.08
Gemini-1.5-Pro 5.40 16.38 27.06 81.13 0.19 14.59 3.79 5.30 16.89 28.91 81.93 0.00 13.76 4.08
Claude-3.5-Sonnet 5.01 15.54 26.32 81.76 0.65 9.39 3.53 5.94 16.65 27.52 81.76 0.46 12.11 4.04
GPT-4V 4.97 14.86 26.62 81.75 0.37 14.17 3.69 5.24 15.65 27.62 82.37 0.20 19.52 4.13
GPT-4o 4.93 15.59 27.02 81.11 0.27 13.20 4.01 5.57 16.36 28.37 81.84 0.36 18.87 4.22
Interleaved Text and Image Data for Pre-training.  MMSci includes full article content and figures, naturally forming interleaved text and image data suitable for pre-training LVLMs (Lin et al., 2023). We discuss the utilization of this interleaved data in Section 7.
6 Benchmark Evaluation Results
Evaluated Models.  We evaluated a range of open-source and proprietary LVLMs, including Kosmos-2 (Peng et al., 2023), Qwen-VL-7B-Chat (Bai et al., 2023a), Qwen2-VL-2B, and Qwen2-VL-7B (Wang et al., 2024a), the LLaVA1.5 and LLaVA-NeXT(1.6) models (Liu et al., 2024, 2023a), IDEFICS2 (Laurençon et al., 2024b) and IDEFICS3 (Laurençon et al., 2024a), the InternVL2 series (Chen et al., 2024b), and Llama3.2-11B-Vision (Team, 2024). For proprietary models, we evaluated Gemini-1.5-Flash and Gemini-1.5-Pro (Reid et al., 2024), Claude-3-Opus (Anthropic, 2024a), Claude-3.5-Sonnet (Anthropic, 2024b), GPT-4V, and GPT-4o (Achiam et al., 2023). The exact model versions used are detailed in Appendix A.2.1.
Table 4: Accuracies (%) of models and human experts on multiple-choice questions. Setting I, II, and III denote Fig2Cap, SubFig2Cap, and SubCap2Fig, respectively.
Model I II III Avg.
@@ 6 Benchmark Evaluation Results
Kosmos2 23.99 23.95 24.33 24.09
LLaVA1.5-7B 32.74 24.31 22.80 26.75
LLaVA1.6-Mistral-7B 34.76 20.38 24.15 26.60
Qwen-VL-7B-Chat 39.56 19.93 27.83 29.23
InternVL2-2B 42.76 33.07 38.42 38.18
InternVL2-8B 52.78 49.60 40.13 47.62
InternVL2-26B 50.59 57.82 71.63 59.81
IDEFICES2-8B 48.65 25.83 21.10 32.21
IDEFICES3-8B-Llama3 50.42 28.43 29.98 36.57
MiniCPM-V-2.6 53.20 58.27 61.67 57.61
Llama3.2-11B-Vision 54.97 45.04 71.18 57.00
Qwen2-VL-2B 60.61 37.62 55.12 51.30
Qwen2-VL-7B 66.16 73.10 79.80 72.87
Qwen2-VL-7B-MMSci 81.48 88.47 92.91 87.48
@@
Gemini-1.5-Flash 54.77 77.84 64.41 65.24
Gemini-1.5-Pro 62.79 81.41 77.16 73.52
Claude-3-Opus 52.19 53.17 60.23 55.13
Claude-3.5-Sonnet 68.77 85.34 87.16 80.18
GPT-4V 60.43 75.07 76.12 70.45
GPT-4o 67.42 87.40 84.65 79.57
Random Guess 25.86 24.63 20.62 23.24
PhD Experts 64.18 71.64 72.72 69.51
@@
Scientific Figure Captioning Results.  As shown in Table 3, grounding captions in article abstracts consistently improves generation quality across all models by providing essential context. While most models struggle to capture the nuanced semantics and style of ground truth captions, our fine-tuned model effectively learned these subtle details from the training data. Although Qwen-VL-7B-Chat achieves high scores on certain metrics, this is primarily due to its tendency to generate concise outputs. The consistently low CIDEr scores across models highlight the distinct challenges of captioning our figures than natural images.
In terms of LLM-based metrics, which evaluate quality beyond semantic nuance, open-source models significantly underperform compared to proprietary models. For G-Eval, which assesses overall similarity to reference captions, proprietary models achieve superior performance. However, on FActScore, which measures precision in describing specific figure components, our fine-tuned model performs best. Nevertheless, all models fall short of satisfactory performance, highlighting the ongoing challenge of precise scientific figure description.
Multi-choice Question Results.  Table 4 presents the multiple-choice question results across three settings. The Figure-to-Caption (Setting I) task, requiring models to identify correct summaries of multi-panel figures (examples in Figure 7, Appendix), proved most challenging. Our fine-tuned model outperformed the strongest proprietary model by over 10%. For SubFig2Cap (Setting II) and SubCap2Fig (Setting III), proprietary models significantly outperformed open-source models, suggesting limitations in identifying nuanced figure content. While some open-source models (LLaVA1.5, LLaVA1.6, Qwen-VL-7B-Chat) performed at random-chance levels, others (MiniCPM-V-2.6, Llama3.2-11B-Vision, Qwen2-VL-7B) demonstrated strong competitiveness. Although Claude-3.5-Sonnet and GPT-4V led among proprietary models, our fine-tuned Qwen2-VL-7B-MMSci achieved the highest overall performance, validating our training data’s effectiveness.
PhD Expert Evaluations.  We organized the dataset into 10 major scientific categories aligned with the Prolific platform [^4] https://www.prolific.com/: Material Science, Chemistry, Physics, Biochemistry, Environment, Climate Sciences, Earth Sciences, Biological Sciences, Biomedical Sciences, and Health and Medicine. For each category, we selected 75 questions (25 per setting) and recruited three PhD-level evaluators with verified degrees in each domain through Prolific, totaling 30 experts. The evaluators provided two assessments: (1) Question Quality Assessment: Experts rated question clarity and domain-knowledge testing effectiveness on a 5-point scale; and (2) Human Expert Performance: Experts answered questions with a suggested time limit of one minute to establish a human performance baseline.
The 30 PhD experts gave an average question quality score of 4.01 (4.09 for Fig2Cap, 4.03 for SubFig2Cap, and 3.91 for SubCap2Fig), where a score of 4 indicates that the question is clear, answerable, and requires an adequate understanding of the scientific content in the figure. This validates the quality of our benchmark questions. The averaged highest-performing expert results for each domain are reported in Table 4. Notably, our fine-tuned model and leading proprietary models achieved performance surpassing PhD-level experts, who were constrained to one minute per question. This superior performance likely reflects the models’ ability to rapidly process dense scientific information across various domains, highlighting both the task’s complexity and the potential of LVLMs as efficient cross-domain scientific assistants. Detailed evaluation procedures and results are provided in Appendix A.2.3.
@@ 7 A Case Study in Material Sciences
Material science as the subject with the most articles and figures in our dataset, is an important and highly interdisciplinary field that requires knowledge from various subjects. Given its significance, we conducted a case study to explore how our dataset could enhance material science knowledge. Previous research has investigated the application of language models to material science tasks (Walker et al., 2021; Rubungo et al., 2023; Miret & Krishnan, 2024). A recent study (Gruver et al., 2024) demonstrated promising results using LLaMA2 (Touvron et al., 2023) for material generation by representing crystal structures as text strings and training the model to generate these structures. However, LLaMA2’s scientific knowledge may be insufficient for fully understanding material generation principles. To address this limitation, we explored continuous pre-training of LLaMA2 using our interleaved scientific article and figure dataset, aiming to improve the model’s performance on stable material generation tasks.
Visual Pre-Training on MMSci We continuously pre-trained the LLaMA2-7B model on our collected interleaved article text and figure images, using data within materials science as well as other eight related subjects in the same Physical Science category. To achieve that, we leverage LLaVA’s architecture (Liu et al., 2024), equipping LLaMA2 with a pre-trained CLIP ViT-L/14-336 (Radford et al., 2021) as the visual encoder and a 2-layer MLP as the projector. During training, we initially kept the LLM frozen and used data from general domains provided by (Liu et al., 2024) to initialize the projector. We then trained the model on the interleaved text and image data from general domains in MMC4 (Zhu et al., 2024) to further develop its image perception abilities, followed by our collected interleaved articles and figures in MMSci to infuse scientific knowledge. In this stage, we tuned both the LLM and the projector, for one epoch. For the resulting multimodal model, we use its LLM part, named LLaMA2-7B-MMSci, for the subsequent material generation.
@@
Fine-tuning for Materials Generation.  Given the LLM, we further fine-tune it for the material generation task as in (Gruver et al., 2024). Specifically, periodic materials are characterized by a unit cell that repeats infinitely in all three dimensions. Each unit cell is specified by its side lengths ($l_{1}$, $l_{2}$, $l_{3}$) and angles ($\theta_{1}$, $\theta_{2}$, $\theta_{3}$). Within this lattice structure, there are $N$ atoms, each identified by an element symbol, $e_{i}$, and a set of 3D coordinates ($x_{i}$, $y_{i}$, $z_{i}$). Tzhe structure of a bulk material $C$ can be represented by:
@@
Following (Xie et al., 2021; Gruver et al., 2024), we use the MP-20 dataset (Jain et al., 2013) of 45,231 stable materials, where successful generation should produce at least metastable crystals. The training data incorporates both conditional generation prompts (single or multiple conditions) and infilling prompts for masked crystal structure strings. Training is limited to one epoch to maintain diversity in generated materials.
@@
GPT-4o with Few-shot Prompting
GPT-4o 5-shot 0.799 0.898 0.280 0.961 5.421 1.017 1.50% -
GPT-4o 10-shot 0.787 0.820 0.654 0.963 3.976 0.917 4.72% 0.09%
Gruver et al. (2024): LLaMA2 with Task-specific Fine-Tuning
@@
Ours: LLaMA2 with Continuous Pre-Training on MMSci plus Task-specific Fine-Tuning
@@
Results.  For evaluation, we perform unconditional generation of potential stable materials using a temperature of 0.7 to sample 10,000 structures (Xie et al., 2021; Gruver et al., 2024). We assess performance across four metrics: validity (adherence to physical constraints), coverage and property metrics (alignment with ground truth distribution), and stability (percentage of samples deemed metastable by M3GNet (Chen & Ong, 2022) and stable by DFT (Hafner, 2008)). As shown in Table 5, GPT-4o fails at this task without specific training. LLaMA2-7B achieves superior results after continuous pre-training on our interleaved articles and figures, followed by multi-task fine-tuning. This model demonstrates the best performance in compositional validity, coverage precision, and the most important metrics metastability, and stability, highlighting the benefit of our data in enhancing the generative model’s acquisition of scientific knowledge.
@@
Ablation Studies.  To understand the factors contributing to LLaMA2-7B-MMSci’s performance, we explored different pre-training data configurations: using only interleaved data from either MMC4 (general interleaved data) or MMSci, using interleaved data from MMC4 combined with text-only data from MMSci, and using no additional pre-training data, followed by the same fine-tuning setup. As shown in Figure 5, the text-only and interleaved data from MMSci achieved the top-2 overall performance when combined with MMC4 which equips the model to effectively read text and interpret images within scientific articles. Using both articles and figures led to better performance than using text-only data from MMSci, highlighting the importance of understanding both figures and content in scientific literature. In contrast, using only general domain data from MMC4 did not result in improvements, and directly training on MMSci even slightly decreased performance in structure validity. This is likely because incorporating visual information can confuse the model if it has not been sufficiently pre-trained with general interleaved data. Overall, continuous pre-training on our data shows the potential to infuse scientific knowledge that enhances downstream tasks.